OpenStudy (lukecrayonz):

(f(x+h)-f(x))/h if f(x)=-2x^2

OpenStudy (campbell_st):

so its 1st principals... f(x) = -2x^2 f(x + h) = -2(x + h)^2 so f(x +h) - f(x) = -2(x + h)^2 - -2x^2 ------------ ------------------ h h there is probably a limit as h ==> 0 or something like that so distribute f(x + h) then collect like terms look for a common factor that cancels fromt he numerator and denominator.

OpenStudy (lukecrayonz):

Why is f(x+h) equal to -2(x+h)^2?

OpenStudy (campbell_st):

f(x) = -2x^2 f(x+h) means substitute (x +h) where ever x occurs its like f(x) = 3x f(2) = 3*2 = 6